OpenStudy (anonymous):

HELP ME PLEASEE ?? The cone shaped cup has a diameter of 3.25 inches and is 3.75 inches tall. The cylindrical cup has a diameter of 4 inches and height of 4.75 inches. The spherical scoop you have chosen has a diameter of 2.75 inches. Which cup should you choose if you want two scoops of ice to best fit in the cup?

OpenStudy (amistre64):

"best fit" is ill defined ... just use the larger one

OpenStudy (anonymous):

the cylinder.

OpenStudy (anonymous):

well. maybe not. one sec.

OpenStudy (anonymous):

Yeah the cylinder.

OpenStudy (anonymous):

you have to show your work for the answer.. @amistre64

OpenStudy (amistre64):

you cant show your work for an ill defined term ....

OpenStudy (anonymous):

nah you got to.

OpenStudy (amistre64):

show your work for which couch is prettier .... the plaid striped one or the one that the dog vomited on ....

OpenStudy (anonymous):

lol.

OpenStudy (anonymous):

Use complete sentences and good grammar to write a justification of your choice. Show all the mathematics you used in your calculations. It says this with the problem so I'm sure.. there some work to this !

OpenStudy (amistre64):

:)

OpenStudy (anonymous):



OpenStudy (amistre64):



OpenStudy (anonymous):

so you don't have to add or multiply nothing in this problem ???

OpenStudy (anonymous):

sure but use common sense.

OpenStudy (amistre64):

let y = -3.75/(3.25/2)x + 3.75 represent the cone let x^2+y^2 = 2.75^2 represent the scoop .....

OpenStudy (anonymous):

do i solve this... and it says to show your work so i know the answer but need some calculations

OpenStudy (amistre64):

and thats spose to be 2.75/2 for the radius ....

OpenStudy (anonymous):

way over my head. sorry i couldnt help.

OpenStudy (primeralph):

@ehmuleeee Use your geometry skills here. She knows lot of geometry.

OpenStudy (anonymous):

@primeralph you're funny

OpenStudy (anonymous):

YEAH! @ehmuleeee

OpenStudy (amistre64):

are we mashing the icecream into the container?

OpenStudy (amistre64):

or is it rock hard and breaks a tooth when you bite it?

OpenStudy (anonymous):

why does that matter ?

OpenStudy (anonymous):

how*

OpenStudy (amistre64):

well, if we are to fit volume to volume .... then mashing the icecream into the cone might make it __"fit better"__

OpenStudy (anonymous):

ok i guess we will mash it then

OpenStudy (amistre64):

so, does 2*4/3*pi r^3 <= pi r^2h/3 ?

OpenStudy (anonymous):

sure

OpenStudy (amistre64):

well, we cant even fit it all in by mashing it flat .... maybe we can fit it in by mashing one and seeing of the other can roll about on top .... is the cone edible?

OpenStudy (anonymous):

sorry, but this is confusing to me.. all this math

OpenStudy (primeralph):

@xoxox123 Welcome to earth.

OpenStudy (anonymous):

yes it probably is.

OpenStudy (anonymous):

@amistre64 can you just tell me an easy way to say the answer is a cylinder

OpenStudy (amistre64):

i thought i did ... i tend to give my easiest advice first

OpenStudy (amistre64):

showing that graphic representation of the scoop and cone does not allow for it to hold 2 scoops (it hardly hold one) indicates that the other one should be choosen

OpenStudy (anonymous):

so the cylinder is the best answer ??

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

id choose it, yes.